# Sensitivity adjustable in-line high-temperature sensor based on metal microwire optical Fabry-Perot interferometer

### v1

@@ III Experimental results and discussion
The metal microwire-based FPI sensor temperature performance investigation setup is depicted in Fig. 3. The broadband light from the white light source was launched into the metal microwire-based FPI sensor through the optical fiber connected to the optical circulator. The reflection spectrum of the FPI was recoupled to the optical circulator and recorded using an optical spectrum analyzer (OSA, Anritsu, MS9740A). The metal microwire-based FPI sensor was mounted on a thermostatic heating table and heated from 50°C to 440°C with an incremental step of 10°C. The temperature response of a microwire-based FPI with a length of approximately 88.5 $\mu m$ (Fig. 1(b)) is presented in Fig. 4. Fig. 4(a) illustrates that the reflection interference spectrum shifts to the short-wavelength side (blue shifting) as the temperature increases from 170°C to 200°C. Due to the limited bandwidth of our white light source, we were only able to monitor the FPI sensor’s temperature response within an approximate range of 150°C. Consequently, we identified three resonance dips of the FPIs as the temperature increased from 50°C to 410°C. The results of these observations are presented in Fig. 4(b), (c) and (d). The results indicated that this sensor’s reflection spectrum resonant dips shift to shorter wavelengths with increasing temperature, and their sensitivity is greater than -0.35 nm/°C.
@@ IV Conclusion
Acknowledgements We wish to thank Prof.Yinhong Liao for his assistance in the sensor’s temperature performance investigation experiments. This work is supported by the Chongqing Ph.D. “Through Train” Scientific Research Project (Grant Number: sl202100000172) and Fundamental Research Funds for the Central Universities (Grant Number: SWU021004).

### v2

@@ III Experimental results and discussion
The metal microwire-based FPI sensor temperature performance investigation setup is depicted in Fig. 3. The broadband light from the white light source was launched into the metal microwire-based FPI sensor through the optical fiber connected to the optical circulator. The reflection spectrum of the FPI was recoupled to the optical circulator and recorded using an optical spectrum analyzer (OSA, Anritsu, MS9740A). The metal microwire-based FPI sensor was mounted on a thermostatic heating table and heated from 50°C to 440°C with an incremental step of 10°C. The temperature response of a microwire-based FPI with a length of approximately 88.5 $\mu m$ (Fig. 1(b)) is presented in Fig. 4. Fig. 4(a) illustrates that the reflection interference spectrum shifts to the short-wavelength side (blue shifting) as the temperature increases from 170°C to 200°C. Due to the limited bandwidth of our white light source, we were only able to monitor the FPI sensor’s temperature response within an approximate range of 150°C. Consequently, we identified three resonance dips of the FPIs as the temperature increased from 50°C to 410°C. The results of these observations are presented in Fig. 4(b), (c) and (d). The results indicated that this sensor’s reflection spectrum resonant dips shift to shorter wavelengths with increasing temperature, and their sensitivity is greater than -0.35 nm/°C. . As illustrated in Fig. 4, the high temperature sensitivity of the metal microwire-based FPI results in a significant shift in the resonant wavelength when monitoring temperatures. Consequently, any fringe valley would shift outside the wavelength range of the spectrometer when attempting to cover a broad temperature range. To overcome this limitation, two techniques can be employed: the first is to stitch several resonant valleys together, as demonstrated by Liu et al.[24], and the second is to directly demodulate the temperature-induced FPI cavity length change, as illustrated in Fig. 5.
@@
In this study, the Cr20Ni80 metal microwire was employed as the temperature-sensitive component of the FPI device. Indeed, a variety of alternative metal microwires may be selected for this purpose, including gold wire, platinum wire, copper wire, aluminum wire, and steel wire. The Cr20Ni80 metal microwire was ultimately selected due to its advantageous properties compared to other metal wires. The Cr20Ni80 wire displays a relatively high thermal expansion coefficient of 18×10-6, although this value is lower than that of lead and aluminum wires. Its melting point is considerably higher than that of lead and aluminum wires. Moreover, the melting point of the metal microwire exceeds the softening point of silica glass fiber, which results in the FPI’s damage temperature threshold extending to the softening point of the silica glass fiber. Furthermore, it is more cost-effective compare to gold wire and platinum wire.
Unlike the silica material, the metal wire may absorb specific directions and wavelengths of external electromagnetic fields, resulting in a thermal effect that affects the accuracy of temperature measurements. However, given the small and short dimensions of the metal wire, the heat generated by our typical exposure to external electromagnetic fields can be considered negligible. Nevertheless, the wire does exhibit a capacity to absorb shorter wavelength electromagnetic fields, including light, infrared, and far infrared radiation. The underlying mechanisms involved are complex and warrant further investigation.
@@ IV Conclusion
Acknowledgements We wish to thank Prof. Yinhong Liao for his assistance in the sensor’s temperature performance investigation experiments.
Founding This work is supported by the Chongqing Ph.D. “Through Train” Scientific Research Project (Grant Number: sl202100000172) and Fundamental Research Funds for the Central Universities (Grant Number: SWU021004).
@@
Author’s Contributions D.Duan designed the study, performed the experiments, analyzed the data, and drafted the manuscript. Z.Zhao participated in the experiments and data analysis. Y.Xie participated in the study design and coordination and helped draft the manuscript. All authors read and approved the final version of the manuscript.